OpenStudy (abmon98):

the parametric equations of a curve are x=t-e^2t y=t+e^2t where take t all real values express dy/dx in terms of t. ii) hence find the value of t for which the gradient of the curve is 3 giving your answers in logarithmic form iii) show that for points on the curve the greatest value of x is 1/2(ln1/2-1)

OpenStudy (abmon98):

i)\[x=t-e ^{2t}\rightarrow dx/dt=1-(e ^{2t}\times d/dt(2t))=1-2e ^{2t}\] \[y=t+e ^{2t}\rightarrow dy/dt=1+(e ^{2t}\times d/dt(2t))=1+2e^2t\] \[dy/dx=dy/dt \div dx/dt=(1+2e^2t) \div (1-2e^2t)\]

OpenStudy (abmon98):

\[3-6e ^{2t}=1+2e ^{2t}\rightarrow 8e ^{2t}=2\rightarrow e ^{2t}=1/4\rightarrow 2t*lne=\ln(1)-\ln(4)\rightarrow t=-\ln(4)/2\]

OpenStudy (kainui):

So far so good. =)

OpenStudy (abmon98):

dy/dx=1+2e^2t/1-2e^2t 0=1+2e^2t -1/2=e^2t -(ln(1)-ln(2))=lne^2t -(-ln(2)=2t t=ln2/2

OpenStudy (abmon98):

x=t-e^2t =ln(2)/2-e^2(ln(2)/2) =ln(2)/2-2

OpenStudy (abmon98):

@eliassaab

OpenStudy (anonymous):

For the first question t=-ln(2)

OpenStudy (abmon98):

t=-ln(4)/2